### v1

# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

Abstract.
We examine various density results related to the solutions of the non-local heat equation at a specific time slice, focusing on two distinct models: one with homogeneous Dirichlet boundary condition and the other with singular boundary data. We explore both the qualitative and quantitative aspects of the approximations. Additionally, we address Calderón-type inverse problems for these parabolic models, where we recover the potentials by analyzing the solutions either on the boundary or at a particular time slice. In both the density results and the Calderón type inverse problems, the Pohozaev identity plays a crucial role. Finally, in the last section, we apply the Pohozaev identity to a specific elliptic eigenvalue problem and demonstrate that the eigenfunctions, when divided by an appropriate power of the distance function, can not vanish on any non-empty open subset of the boundary.
1. Introduction
In this article, we address various qualitative and quantitative approximation properties of solutions to certain non-local heat equation. Specifically, we consider two equations where one has homogeneous Dirichlet boundary condition and the other has singular boundary data. For the singular boundary data, we analyze the approximation properties by deriving an observability estimate through the Pohozaev identity [see [ROS14b]]. Using the same Pohozaev identity and density results, we also recover the potentials for both the systems by analyzing the solutions either at a specific time slice or on the boundary. We consider the following two equations:
(1) $\partial_{t}u_{f}+(-\Delta)^{a}u_{f}+qu_{f}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{f}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ u_{f}(0,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
and
(2) $\partial_{t}u_{F}+(-\Delta)^{a}u_{F}+qu_{F}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{F}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{u_{F}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ u_{F}(0,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega.\end{aligned}\right.$
Here $\Omega\subset\mathbb{R}^{N}$ is a bounded $C^{1,1}$ domain with boundary $\partial\Omega$ and $a\in(0,1)$. The functions $u_{f},u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ are the unknowns and the potential $q\in C_{c}^{\infty}\left((0,T)\times\Omega\right)$. Furthermore, for (1), we assume the initial condition $f\in C_{c}^{\infty}(\Omega)$ and for (2) we assume the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. The distance function ${d:\Omega\to\mathbb{R}_{\geq 0}}$ is the following:
$d(x):=\inf\limits_{y\in\partial\Omega}\|x-y\|_{2},$
where $\|\cdot\|_{2}$ is the classical Euclidean norm in $\mathbb{R}^{N}$. To ensure uniqueness, we choose $q$ in such a way that zero is not an eigenvalue for either equation (1) or (2), corresponding to the cases where $f\equiv 0$ or $F\equiv 0$, respectively. Regarding approximation, in [RS17], the authors investigate a qualitative and quantitative approximation properties of solutions to a specific non-local heat equation, where they vary the $N+1$ Dirichlet data in the exterior domain. They showed that the solution space is dense in $\mathrm{L}^{2}$ in the interior by leveraging the unique continuation property: if a function $u$ satisfies $u=(-\Delta)^{a}u=0$ in some open set $W\subset\mathbb{R}^{N}$, then it must vanish identically, i.e., $u\equiv 0$ [GSU20]. The precise result is as follows:
Theorem 1 ([RS17]).
Let $a\in(0,1)$ and $B_{1}$ is the unit ball in $\mathbb{R}^{N}$ centered at origin. Let $W\subset\mathbb{R}^{N}$ be a non-empty open set such that $\overline{B}_{1}\cap\overline{W}$ is empty. Let for $f\in C_{c}^{\infty}((-1,1)\times W)$, the function $u_{f}:((-1,1)\times\mathbb{R}^{N})\to\mathbb{R}$ be the solution of the following equation:
(3) $\partial_{t}u_{f}+(-\Delta)^{a}u_{f}=&0\qquad\qquad\mbox{in}\ (-1,1)\times B_{1}\\ u_{f}=f&\qquad\qquad\ \mbox{in}\ (-1,1)\times\mathbb{R}^{n}\setminus B_{1}\\ u_{f}(-1,\cdot)=&f\qquad\qquad\mbox{in}\ B_{1}.\end{aligned}\right.$
Then the solution set
[…]
is dense in $\mathrm{L}^{2}((-1,1)\times B_{1})$.
The quantitative analysis enables the authors to derive an estimate for the cost of approximation, as described in [RS17]. This requires the use of variational techniques, as presented in [RS17], [FCZ00], and [Lio92]. A qualitative version of similar approximation properties for the non-local elliptic model is presented in [GSU20], where the authors investigate the Calderón type problem for the fractional heat equation. A comparable result in the $C^{k}$ space is established in [DSV19], where the authors demonstrate that any $C^{k}$ function can be approximated by smooth solutions to the non-local heat equation, with the approximation property governed by the fractional component of the equation only. However, in this article, we vary either the $N$ dimensional initial data or boundary data, and obtain $\mathrm{L}^{2}$ approximation on an $N$-dimensional interior slice. In particular, we vary the initial data $f\in C_{c}^{\infty}(\Omega)$ for the equation with homogeneous Dirichlet boundary data (1), and the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ for the equation (2). We demonstrate that for a fix time slice $T>0$, the solution sets ${\{u_{f}:f\in C_{c}^{\infty}(\Omega)\}}$ and ${\{u_{F}:F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ corresponding to the equations (1) and (2) respectively, are dense in $\mathrm{L}^{2}(\Omega)$. We think such result and techniques executed here do not depend on $a\in(0,1)$, therefore this could be extended to the classical heat equation i.e. $a=1$.
These results are partially motivated by the stability results for Calderón type inverse problems for the non-local heat equation, as studied in [RS17]. Our work can be seen as a continuation of the contributions made in [RS17] and [FCZ00], extending the analysis to boundary controllability in the presence of singular boundary data. Additionally, the study in [ROS14b] serves as a key motivation for exploring boundary controllability and Calderón-type inverse problems in the context of both singular boundary and initial data problem.
The density argument for the case of the singular boundary data (2) requires observability estimates, which we derive using the Pohozaev estimate, as demonstrated in [ROS14b]. Using density results and the same arguments presented in the proof of observability estimate, we solve the Calderón type inverse problems for both the non-local heat equations (1) and (2), where we analyze the solutions at a particular time slice $T>0$ for (1) and where we analyze the solutions on the boundary for (2). The study of the Calderón type inverse problem for the non-local elliptic equation can be found in [GLX17], [GRSU20], [Sal17], [Gho21]. The main results in this article are the following:
Theorem 2.
Let $f\in C_{c}^{\infty}(\Omega)$ and let $u_{f}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (1). Then, for a fix time $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$.
Theorem 3.
Let $F\in C_{c}^{\infty}((0,T)\times\Omega)$ and let $u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (2). Then, for a fix time $\tilde{T}>0$, the set
$\left\{u_{F}(\tilde{T},\cdot):F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\right\}$
is dense in $L^{2}(\Omega)$.
Theorem 4.
Let $a\in(0,1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a}(\Omega))$ satisfy: for $i=1,2$
(4) $\partial_{t}u_{i}+(-\Delta)^{a}u_{i}+q_{i}u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ u_{i}(0,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $q_{1},q_{2}\in C_{c}^{\infty}((0,T)\times\Omega)$ and $f\in C^{\infty}_{c}(\Omega)$. Furthermore, $q_{1},q_{2}$ be such that zero isn’t an eigenvalue for the same equation with zero initial data. If
$\frac{u_{1}}{d^{a}}(t,x)=\frac{u_{2}}{d^{a}}(t,x),\quad\forall\,(t,x)\in(0,T]\times\partial\Omega\ \mbox{and}\ \forall\,f\in C_{c}^{\infty}(\Omega),$
then
$q_{1}\equiv q_{2}.$
Theorem 5.
Let $a\in(0,1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a-1(2a)}(\Omega))$ satisfy: for $i=1,2$
(5) $\partial_{t}u_{i}+(-\Delta)^{a}u_{i}+q_{i}u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{u_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ u_{i}(0,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $q_{1},q_{2}\in C_{c}^{\infty}((0,T)\times\Omega)$ and $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. Furthermore, $q_{1},q_{2}$ be such that zero isn’t an eigenvalue for (4) with zero boundary data. If
$u_{1}(T,\cdot)\equiv u_{2}(T,\cdot)\qquad\forall\,F\in C_{c}^{\infty}((0,T)\times\Omega)$
then
$q_{1}\equiv q_{2}.$
In fact, we will show statements of theorem 4 and theorem 5 are equivalent. Next we describe solution spaces, well posedness of the solutions and various regularity results. Many of the cases, the notion of the solutions for non-local heat equation can be extended to non-local heat equation as the perturbation potential $q$ lies in $C_{c}^{\infty}((0,T)\times\Omega)$ [Paz12], [BJ12].
1.1. Preliminaries: Fractional Laplacian and fractional Sobolev space
Let $0\leq a\leq 1$. The fractional Laplacian operator $(-\Delta)^{a}$ is defined over the space of Schwartz class functions $\mathcal{S}(\mathbb{R}^{N})$ as
$(-\Delta)^{a}u(x):=\mathcal{F}^{-1}\left\{|\xi|^{2a}\hat{u}(\xi)\right\},\quad\forall\,x\in\mathbb{R}^{N},$
where $\hat{.}$ and $\mathcal{F}^{-1}$ denotes the Fourier and inverse Fourier transformation respectively. There are many equivalent definitions of fractional Laplacian [see [Kwa17]]. One of such equivalent definition we use throughout this article is given by the Cauchy principle value: for $0\leq a\leq 1$
$(-\Delta)^{a}u(x)=C_{n,a}\ \mbox{p.v}\ \int_{\mathbb{R}^{N}}\frac{u(x)-u(y)}{|x-y|^{N+2a}}\,\rm{d}y,\qquad\forall\,x\in\mathbb{R}^{N},$
where ${C_{n,a}=\frac{4^{a}\Gamma\left(\frac{n}{2}+a\right)}{\pi^{\frac{n}{2}}\Gamma(-a)}}$. Throughout the article, $\Gamma$ stands for the usual Gamma function. We define the following Sobolev space: for $s\in(0,1)$
$H^{s}(\mathbb{R}^{N}):=\{u\in\mathbb{S}^{\prime}(\mathbb{R}^{N}):\langle\xi\rangle^{s}\hat{u}\in\mathrm{L}^{2}(\mathbb{R}^{N}).\},$
where ${\langle\xi\rangle=(1+\|\xi\|_{2}^{2})^{\frac{1}{2}}}$ and $\mathbb{S}^{\prime}(\mathbb{R}^{N})$ is the space of tempered distributions in $\mathbb{R}^{N}$, equipped with the norm
$\|u\|_{H^{s}(\mathbb{R}^{N})}=\|\langle\xi\rangle^{s}\hat{u}\|_{\mathrm{L}^{2}(\mathbb{R}^{N})}.$
The fractional Laplacian extends as a bounded linear map
$(-\Delta)^{a}:H^{s}(\mathbb{R}^{N})\to H^{s-2a}(\mathbb{R}^{N}).$
We also introduce few spaces that are essential for describing the solution spaces and various regularity results throughout the article.
$H^{s}(\Omega):=$$\{u=v|_{\Omega}:v\in H^{s}(\mathbb{R}^{N})\}.$
$C^{s}(\Omega):=$$\{u\in C(\Omega):|u(x)-u(y)|\leq M|x-y|^{s},\,\forall\,x,y\in\Omega,M\geq 0\}$
$C^{s}_{c}(\mathbb{R}^{N}):=$$\{u\in C^{s}(\mathbb{R}^{N}):\,\mbox{supp}\,u\ \mbox{compactly contained in}\ \mathbb{R}^{N}\}.$
where $\Omega$ is a bounded $C^{1,1}$ domain and $s\in(0,1)$. We denote $v|_{\Omega}$ as the restriction of the function $v$ in $\Omega$. $H^{s}(\Omega)$ is a Sobolev space equipped with the norm
$\|u\|_{H^{s}(\Omega)}=\|u\|_{H^{s}(\Omega)}=\|\langle\xi\rangle^{s}\hat{u}\|_{\mathrm{L}^{2}(\Omega)}.$
The definition of the space $C^{s}(\Omega)$ can be extended to the set when $\Omega$ is closed i.e., $\Omega=\overline{\Omega}$ or $\Omega=\mathbb{R}^{N}$. We define the following seminorm associated with the space $C^{s}(\Omega)$:
$[u]_{C^{s}(\Omega)}:=\inf_{M\geq 0}\left\{M:|u(x)-u(y)|\leq M|x-y|^{s},\,\forall\,x,y\in\Omega\right\}.$
The spaces $C_{c}^{s}(\mathbb{R}^{N})$ and $C^{s}(\overline{\Omega})$ $(\Omega$ is bounded)))) can be endowed with the following norms respectively
$\|u\|_{C^{s}_{c}(\mathbb{R}^{N})}:=$ $\sup\limits_{x\in\mathbb{R}^{N}}|u(x)|+[u]_{C^{s}(\mathbb{R}^{N})},$
$\|u\|_{C^{s}(\overline{\Omega})}:=$ $\sup\limits_{x\in\overline{\Omega}}|u(x)|+[u]_{C^{s}(\Omega)}.$
Note that $C^{s}(\overline{\Omega})$ is a Banach space with respect to the norm described abouve.
The definition of $C^{s}(\Omega)$ ((((or $C_{c}^{s}(\mathbb{R}^{N}))$ can be extended to any $\beta>0$. We can express $\beta=k+s$, where $k\in\mathbb{N}\cup\{0\}$ and $s\in(0,1)$. The space $C^{\beta}(\Omega)$ is the following
$C^{\beta}(\Omega):=\left\{D^{\alpha}u\in C(\Omega):|D^{k}u(x)-D^{k}u(y)|\leq M|x-y|^{s},\,\forall\,x,y\in\Omega,M\geq 0\right\}$
where ${\alpha=(\alpha_{1},\cdots,\alpha_{N})\in(\mathbb{N}\cup\{0\})^{N}}$ and ${\sum\limits_{i=1}^{N}\alpha_{i}\leq k}$ and $D^{\alpha}$ is the classical derivative of order $\alpha$.
1.2. Well posedness of the solution
Let us define the following space: for $s\in(0,1)$
[…]
The wellposedness of the two equations has been studied in [CGCV22]. It has been shown that under the assumption $f\in\mathrm{L}^{1}(\Omega,d^{a})$ there exists a solution $u_{f}\in\mathrm{L}^{1}((0,T);\mathrm{L}^{1}(\Omega,d^{a}))$ to (1). In the same article [CGCV22], it has been shown that under the assumption $F\in\mathrm{L}^{1}((0,T)\times\partial\Omega)$ there exists a solution $u_{f}\in\mathrm{L}^{1}((0,T);\mathrm{L}^{1}(\Omega,d^{a}))$. We choose $q$ such that zero is not an eigenvalue for both the equations (1), (2) i.e., if $u_{0}\in\mathrm{L}^{1}((0,T);\mathrm{L}^{1}(\Omega,d^{a}))$ satisfies either (1) or (2), corresponding to $f\equiv 0$ or $F\equiv 0$ respectively, then $u_{0}\equiv 0$. This ensures the uniqueness of the solutions also.
Let $P(t,x,y)$ be the fractional heat semigroup corresponding to the operator $\partial_{t}+(-\Delta)^{a}+q$ with homogeneous Dirichlet boundary condition and $\delta_{x}$(Dirac delta) as initial condition. Then the quantity ${\lim\limits_{\begin{subarray}{c}y\to w\\ w\in\partial\Omega\end{subarray}}\frac{P(t,x,y)}{d^{a-1}(y)}}$ is a well defined quatity [see [CGCV22]]. Well defineness of that particular quantity comes from existence of the same limit function for the corresponding elliptic Green function with homogeneous Dirichlet boundary condition [see [AGCV23]]. The solution $u_{f}$ corresponding to (1) or $u_{F}$ corresponding to (2) can be uniquely expressed as
[…]
$u_{F}(t,x)=\int_{0}^{t}\int_{\partial\Omega}\lim\limits_{\begin{subarray}{c}y\to w\\ w\in\partial\Omega\end{subarray}}\frac{S(t-s,x,y)}{d^{a-1}(y)}h(s,w)$
We can further show if $f\in\mathrm{L}^{2}(\Omega)$, then the solution to (1) belongs to the following space
$u_{f}\in\mathrm{L}^{2}((0,T);H^{a}(\mathbb{R}^{N})),$
given by the same expression above [see [CGCV22], [Gru17]]. The smoothness of the fractional heat semigroup for $t>0$ and away from boundary $\partial\Omega$ [FRRO16], [DZZ23] implies $u_{f}\in C^{\infty}\left((0,T)\times\Omega\right)$. The fractional heat semigroup can be expressed in the following way:
$P(t,x,y)=\sum\limits_{n\in\mathbb{N}}e^{-t\lambda_{n}}\phi_{n}(x)\phi_{n}(y)$
where $\lambda_{1}<\lambda_{2}\leq\dots\leq\lambda_{n}\leq\dots+\infty$ are the eigenvalues and $\phi_{n}$’s are the eigenfunctions of the following Dirichlet valued problem
(6) $\left\{\begin{array}[]{ll}((-\Delta)^{a}+q)\phi_{n}=\lambda_{n}\phi_{n}&\mbox{in}\ \Omega\\ \phi_{n}=0&\mbox{in}\ \mathbb{R}^{N}\setminus\Omega.\end{array}\right.$
Various interior regularity results corresponding to (1) can be found in [ROV18], [FRRO16]. In [FRRO16], authors have proved the following results corresponding to (1):
  • ((((a))))
    For eact $t_{0}>0$
    (7) $\sup\limits_{t\geq t_{0}}\|u_{f}(t,\cdot)\|_{C_{c}^{a}(\mathbb{R}^{N})}\leq C_{1}(t_{0})\|f\|_{\mathrm{L}^{2}(\Omega)}.$
  • ((((b))))
    For eact $t_{0}>0$
    (8) $\sup\limits_{t\geq t_{0}}\left\|\frac{u_{f}(t,\cdot)}{d^{a}}\right\|_{C^{a-\epsilon}\left(\overline{\Omega}\right)}\leq C_{1}(t_{0})\|f\|_{\mathrm{L}^{2}(\Omega)},$
  • ((((c))))
    for each $t_{0}>0$
    (9) $\sup\limits_{t\geq t_{0}}\left\|\frac{d^{j}}{dt^{j}}u_{f}(t,\cdot)\right\|_{C^{a}_{c}(\mathbb{R}^{N})}\leq C_{1,j}(t_{0})\|f\|_{\mathrm{L}^{2}(\Omega)}$
for any $\epsilon\in(0,a)$. Here $C_{1}(t_{0})$,$C_{1}^{j}(t_{0})$ are positive constants. Authors used various eigenfunction estimates corresponding to (6) [see [Gru15b], [Gru15a], [AGCV23]]. The following interior regularity results can be found in [FRRO16].
Theorem 6 ([FRRO16]).
Let $u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a}(\Omega)\right)$ be the solution of (1) with $f\in\mathrm{L}^{2}(\Omega)$. Then the following holds
  • $\bullet$:
    For $t>0$, $u(t,\cdot)\in C^{a}(\mathbb{R}^{N})$ and, for every $\beta\in[a,1+2a)$, $u(t,\cdot)$ is of class $C^{\beta}(\Omega)$ and
    (10) $[u(t,\cdot)]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
  • $\bullet$:
    The function ${\frac{u(t,\cdot)}{d^{a}}}\Big{|}_{\Omega}$ can be continuously extended to $\overline{\Omega}$. Moreover, there exists $\alpha\in(0,1)$ such that ${\frac{u}{d^{a}}\in C^{\alpha}(\overline{\Omega})}$. In addition, for $\beta\in[\alpha,a+\alpha]$, the following estimate holds
    (11) $\left[\frac{u}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
Here $\Omega_{\delta}:=\{x\in\Omega:d(x,\partial\Omega)\geq\delta\}$ and $[\cdot]_{C^{\beta}(\Omega_{\delta})}$ denotes the Hölder seminorm:
$\left[\frac{u(t,x)}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}=\sup\limits_{x,y\in C^{\beta}(\Omega_{\delta}),x\neq y}\frac{|D^{\lfloor\beta\rfloor}u(t,x)-D^{\lfloor\beta\rfloor}u(t,y)|}{|x-y|^{\beta-\lfloor\beta\rfloor}}$
where $\lfloor\beta\rfloor$, denotes the greatest integer function i.e $\lfloor\beta\rfloor:=\sup\{z\in\mathbb{Z}:z\leq\beta\}$.
With (7), (8), (10), (11) in hand, the solution $u_{f}$ corresponds to (1) satisfies the fractional version of Pohozaev identity. For all $t>0$:
(12) $\int_{\Omega}(x\cdot\nabla u)(-\Delta)^{a}u\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $u(-\Delta)^{a}u\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma,$
where $\nu$ is the unit outward normal to $\partial\Omega$ at the point $x$ and $\Gamma$ is the gamma function. Proof of this statement is established in the articles [ROS14b], [FRRO16].
Next we introduce $\mu$ transmission spaces.
1.3. $\mu$ transmission spaces
All the notations and definition here are well explained in [Gru14], [Gru15a], [Gru16b], [Gru16a], [Gru18], [Gru20]. We only recall some important definitions and concepts here.
[…]
Let us denote a function $\tilde{d}$ which is of the form $d(x):=\text{dist}(x,\partial\Omega)$ near boundary and extended smoothly into a positive function in $\Omega$. We define the space
$\mathcal{E}_{\mu}\left(\overline{\Omega}\right):=e^{+}\left\{u(x)=\tilde{d}^{\mu}v(x):v\in C^{\infty}\left(\overline{\Omega}\right)\right\},$
for Re$(\mu)>-1$. For general $\mu$, we refer readers to [Gru15a], [See71]. Next we define pseudodifferential operator ($\psi$do) $\mathcal{P}$ on $\mathbb{R}^{N}$, which is defined from a symbol $p(x,\xi)$ on ${\mathbb{R}^{N}\times\mathbb{R}^{N}}$ in the following way:
$\mathcal{P}u=p(x,D)u=\text{Op}((p(x,\xi))u=(2\pi)^{-n}\int e^{ix\dot{\xi}}p(x,\xi)\hat{u}\,\rm{d}\xi=\mathcal{F}_{\xi\to x}^{-1}((p(x,\xi)\hat{u}(\xi)),$
where $\hat{u}$ is the Fourier transformation of $u$. For the theory of calculus, we refer to books and notes like [Hor63], [Hör66], [Tay06], [Gru08]. We define the symbol space $S^{m}_{1,0}\left(\mathbb{R}^{N}\times\mathbb{R}^{N}\right)$ as follows:
$S^{m}_{1,0}:=\left\{\begin{aligned} p(x,\xi)\in C^{\infty}\left(\mathbb{R}^{N}\times\mathbb{R}^{N}\right):&|\partial_{x}^{\beta}\partial_{x}^{\alpha}p(x,\xi)|\sim O\left(\langle\xi\rangle^{m-|\alpha|}\right),\\ &\forall\alpha,\beta\in\mathbb{N}\cup\{0\},\ \mbox{for some}\ m\in\mathbb{C},\end{aligned}\right.$
where ${\langle\xi\rangle:=\left(1+\|\xi\|^{2}_{2}\right)^{\frac{1}{2}}}$ and $O$ is the big-O function captures the order. The exponent $m$ is called the order of $\mathcal{P}($and $p)$. Note that the operator $\left((-\Delta)^{a}+q\right)$ is a $\psi$do operator of order 2a. Depending on the context, the operator $\mathcal{P}_{+}=r^{+}\mathcal{P}e^{+}$ denotes the truncation of $\mathcal{P}$ to +$\mathbb{R}^{N}_{+}$ or to $\Omega$. We define the following $\mathrm{L}^{2}$ spaces:
$H^{s}\left(\mathbb{R}^{N}\right):=\left\{u\in\mathcal{S}^{\prime}\left(\mathbb{R}^{N}\right)|\mathcal{F}^{-1}\left(\langle\xi\rangle^{s}\hat{u}\right)\in\mathrm{L}^{2}\left(\mathbb{R}^{N}\right)\right\},$
$\dot{H}^{s}\left(\overline{\Omega}\right):=\left\{u\in H^{s}\left(\mathbb{R}^{N}\right)|\ \mbox{supp}\ u\subset\overline{\Omega}\right\},$
$\overline{H}^{s}(\Omega):=\left\{u\in\mathcal{D}^{\prime}(\Omega)|u=r^{+}U\ \mbox{for a}\ U\in H^{s}\left(\mathbb{R}^{N}\right)\right\}.$
Next we define order reducing operator on $\mathbb{R}^{N}$: for $t\in\mathbb{R}$
$\Xi^{t}_{\pm}:=Op\left(\chi^{t}_{\pm}\right),\quad\chi^{t}_{\pm}:=\left(\langle\xi^{\prime}\rangle\pm i\xi_{n}\right)^{t}\ \mbox{where}\ \xi=(\xi^{\prime},\xi_{n}),\xi^{\prime}:=(\xi_{1},\dots,\xi_{n-1}).$
[…]
$\Xi^{t}_{\pm}:H^{s}\left(\mathbb{R}^{N}\right)\xrightarrow{\sim}H^{s-t}\left(\mathbb{R}^{N}\right),\ \Lambda^{t}_{\pm}:H^{s}\left(\mathbb{R}^{N}\right)\xrightarrow{\sim}H^{s-t}\left(\mathbb{R}^{N}\right),$
$\Xi^{t}_{\pm}:\dot{H}^{s}\left(\overline{\mathbb{R}^{N}_{\pm}}\right)\xrightarrow{\sim}\dot{H}^{s-t}\left(\overline{\mathbb{R}^{N}_{\pm}}\right),\ r^{+}\Xi^{t}_{\pm}e^{+}:\overline{H}^{s}\left(\overline{\mathbb{R}^{N}_{\pm}}\right)\xrightarrow{\sim}\overline{H}^{s-t}\left(\overline{\mathbb{R}^{N}_{\pm}}\right),$
$\Lambda^{t}_{\pm}:\dot{H}^{s}\left(\overline{\Omega}\right)\xrightarrow{\sim}\dot{H}^{s-t}\left(\overline{\Omega}\right),\ r^{+}\Lambda^{t}_{\pm}e^{+}:\overline{H}^{s}\left(\overline{\Omega}\right)\xrightarrow{\sim}\overline{H}^{s-t}\left(\overline{\Omega}\right).$
We recall $\mu$-transmission spaces with Re $\mu>-1$ as introduced by Hörmander [Hor63], [Hör66] and redefined in [Gru08]:
$H^{\mu(s)}\left(\overline{\mathbb{R}_{+}^{N}}\right):=$ $\Xi_{+}^{-\mu}e^{+}\overline{H}^{s-\mu}\left(\mathbb{R}_{+}^{N}\right),\quad s>\mu-\frac{1}{2}$
$H^{\mu(s)}\left(\overline{\Omega}\right):=$ $\Lambda_{+}^{(-\mu)}e^{+}\overline{H}^{s-\mu}(\Omega),\ \ \quad s>\mu-\frac{1}{2}.$
The space ${{\mathcal{E}_{\mu}\left(\overline{\Omega}\right)}}$ is dense in ${H^{\mu(s)}\left(\overline{\Omega}\right)}$ i.e. [see [Hor63], [Hör66], [Gru08]]
$\cap_{s}H^{\mu(s)}\left(\overline{\Omega}\right)=\mathcal{E}_{\mu}\left(\overline{\Omega}\right).$
Also we have the following characterization [see [Gru08]]:
$H^{\mu(s)}\left(\overline{\Omega}\right)\left\{\begin{aligned} &=\dot{H}^{s}\left(\overline{\Omega}\right)\ \ \ \mbox{if}\ |s-\mu|<\frac{1}{2},\\ &\subset\dot{H}^{s}\left(\overline{\Omega}\right)+e^{+}\tilde{d}^{\mu}\overline{H}^{s-\mu}(\Omega),\ \ \mbox{if}\ s>\mu+\frac{1}{2},s-\mu-\frac{1}{2}\not\in\mathbb{N}.\end{aligned}\right.$
We now state some well posedness and regularity results corresponding to (1) and (2) in $\mu$-transmission spaces. These results are useful to obtain an integration by parts formula. Proof of these non-local parabolic well posedness and regularity results can be found in [Gru23], [FRRO16].
Theorem 7.
Let $f\in C_{c}^{\infty}(\Omega)$, then the solution $u_{f}$ to (1) has the following regularity:
$u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right).$
It follows from [Gru23]. The function $v=u_{f}-f$ satisfies
$\partial_{t}v+(-\Delta)^{a}v+qv=&(-\Delta)^{a}f+qf\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ v=&0\qquad\qquad\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ v(0,\cdot)=&0\qquad\qquad\qquad\qquad\qquad\mbox{in}\ \Omega.\end{aligned}\right.$
As $f\in C_{c}^{\infty}(\Omega)$, we have $(-\Delta)^{a}f+qf\in\mathrm{L}^{2}(\Omega)$. Hence $v\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right)$ [Gru23]. Furthermore, as $f\in\mathcal{E}_{a}(\overline{\Omega})$, we have
$u_{f}\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right).$
Theorem 8 ([Gru23]).
Let $F\in\mathrm{L}^{2}\left((0,T);H^{a+\frac{1}{2}}(\partial\Omega)\right)\cap\overline{H}^{1}\left((0,T);H^{\epsilon}(\partial\Omega)\right)$, for some $\epsilon>0$ and let $F(0,\cdot)=0$. Then the solution $u_{F}$ to (2) has the following regularity:
$u_{F}\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)\cap\overline{H}^{1}\left((0,T);\mathrm{L}^{2}(\Omega)\right).$
All the above results are also true for the solution of the adjoint system corresponding to (1), given by:
(13) $-\partial_{t}u_{f^{*}}+(-\Delta)^{a}u_{f^{*}}+qu_{f^{*}}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ u_{f^{*}}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ u_{f^{*}}(T,\cdot)=&f^{*}\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $f^{*}\in\mathrm{L}^{2}(\Omega)$. It just comes from change of variable $t\to T-t$. We like to emphasize that Pohozaev identity (12) holds for $u_{f^{*}}$ also.
With all these regularity results in hand we are going to state integration by parts formula: If ${u,v\in H^{(a-1)(s)}(\overline{\Omega})}$ with ${s>a-\frac{1}{2}}$, then
$\int_{\Omega}v(-\Delta)^{a}u-u(-\Delta)^{a}v=\Gamma(a)\Gamma(a+1)\int_{\partial\Omega}\partial_{\nu}\left(\frac{u}{d^{a-1}}\right)\frac{v}{d^{a-1}}-\partial_{\nu}\left(\frac{v}{d^{a-1}}\right)\frac{u}{d^{a-1}},$
where ${\partial_{\nu}}$ is the normal derivative at the boundary. Furthermore, if $u\in H^{(a-1)(s)}(\overline{\Omega})$ and $v\in H^{a(s)}(\overline{\Omega})$, then
(14) $\int_{\Omega}v(-\Delta)^{a}u-u(-\Delta)^{a}v=-\Gamma(a)\Gamma(a+1)\int_{\Omega}\frac{u}{d^{a-1}}\frac{v}{d^{a}}.$
Here the boundary contributions are completely local. The formula was first proved for fractional laplacian in [Aba13] and later generalized for any classical pseudodifferential operator of order $2a$ in [Gru16a], [Gru20]. We would like to note that for any $t>0$, the integration by parts formula holds for the functions $u_{f}$ and $u_{F}$, which are the solutions to (1) and (2), respectively, where $f\in\mathrm{L}^{2}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\partial\Omega)$. Thanks to theorem 7 and theorem 8, the formula clearly holds for $f\in C_{c}^{\infty}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\Omega)$. Let $f\in\mathrm{L}^{2}(\Omega)$, then there exists a sequence $f_{m}\in C_{c}^{\infty}(\Omega)$ such that $f_{m}\to f$ in $\mathrm{L}^{2}(\Omega)$. For all $m\in\mathbb{N}$, we have
(15) $\int_{\Omega}u_{f_{m}}(t,\cdot)(-\Delta)^{a}u_{F}(t,\cdot)-$ $u_{F}(t,\cdot)(-\Delta)^{a}u_{f_{m}}(t,\cdot)$
$=$ $-\Gamma(a)\Gamma(a+1)\int_{\Omega}\frac{u_{F}}{d^{a-1}}(t,\cdot)\frac{u_{f_{m}}}{d^{a}}(t,\cdot).$
Thanks to [CGCV22], we can represent the solution in the following way:
$u_{f_{m}}(t,\cdot)=\sum\limits_{n\in\mathbb{N}}e^{-t\lambda_{n}}\phi_{n}(x)\langle\phi_{n},f_{m}\rangle,$
where $\phi_{n}$’s are the eigenfunctions corresponding to eigenvalue $\lambda_{n}$, satisfying (6). For a given $t>0$, we have $u_{f_{m}}(t,\cdot)\to u_{f}(t,\cdot)$, $(-\Delta)^{a}u_{f_{m}}(t,\cdot)\to(-\Delta)^{a}u_{f}(t,\cdot)$ and ${\frac{u_{f_{m}}}{d^{a}}(t,\cdot)\to\frac{u_{f}}{d^{a}}(t,\cdot)}$ in $\mathrm{L}^{2}(\Omega)$. Hence, taking $m\to+\infty$ in (15), we obtain
(16) $\int_{\Omega}u_{f}(t,\cdot)(-\Delta)^{a}u_{F}(t,\cdot)-$ $u_{F}(t,\cdot)(-\Delta)^{a}u_{f}(t,\cdot)$
$=$ $-\Gamma(a)\Gamma(a+1)\int_{\Omega}\frac{u_{F}}{d^{a-1}}(t,\cdot)\frac{u_{f}}{d^{a}}(t,\cdot),$
for any $t\in(0,T]$.
We also mention a particular integration by parts formula which we will use throughout the article. Let $w_{1},w_{2}\in H^{a}(\mathbb{R}^{N})$ with $w_{1}\equiv w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$, then
(17) $\int_{\Omega}w_{2}\,(-\Delta)^{a}w_{1}\,\rm{d}x=\int_{\Omega}w_{1}\,(-\Delta)^{a}w_{2}\,\rm{d}x.$
This can be established easily by considering the integral in Fourier space.
Next, we proceed to establish our density results. We present two different proofs for each of these results, addressing both qualitative and quantitative aspects.
2. Density Results:
Proof of theorem 2:
Proof.
We will provide two different proofs. One using integration by parts (14) and another using a variational approach.
First proof: We select the time slice $t=\tilde{T}$. Let the set ${\left\{u_{f}(\tilde{T},\cdot),f\in C_{c}^{\infty}(\Omega)\right\}}$ is not dense in $\mathrm{L}^{2}(\Omega)$. Then by Hahn-Banach theorem there exists $f^{*}\in\mathrm{L}^{2}(\Omega)$ such that $\|f^{*}\|_{L^{2}(\Omega)}=1$ and
(18) $\langle u_{f}(\tilde{T},\cdot),f^{*}\rangle=0,\qquad\forall f\in C_{c}^{\infty}(\Omega).$
Let $u_{f^{*}}\in\mathrm{L^{2}((0,\tilde{T});H^{a}(\Omega)})$ be the solution of the adjoint system (13) with $f^{*}$ as initial data. We use the integration by parts formula (17)
$\int_{\Omega}u_{f}(-\Delta)^{a}u_{f^{*}}=\int_{\Omega}u_{f^{*}}(-\Delta)^{a}u_{f}$
implies
$\int_{\Omega}\partial_{t}u_{f^{*}}u_{f}+u_{f^{*}}\partial_{t}u_{f}=0=\int_{\Omega}\partial_{t}\left(u_{f^{*}}u_{f}\right).$
Integrating with respect to $t$ yields:
$\int_{\Omega}u_{f^{*}}(\tilde{T},\cdot)u_{f}(\tilde{T},\cdot)=\int_{\Omega}u_{f^{*}}(0,\cdot)u_{f}(0,\cdot).$
Thanks to (18) and since $f\in C_{c}^{\infty}(\Omega)$ is arbitrary, we deduce that
$u_{f^{*}}(0,\cdot)\equiv 0.$
Furthermore, the solution can be represented in the following form [see [CGCV22]]:
$u_{f^{*}}(t,x)=\sum\limits_{n\in\mathbb{N}}e^{-\lambda_{n}(\tilde{T}-t)}\phi_{n}(x)\langle\phi_{n},f^{*}\rangle$
where for all $n\in\mathbb{N}$, $\phi_{n}$’s are the eigenfunctions corresponding to (6). We have that $\lambda_{1}<\lambda_{2}\leq\dots\lambda_{n}\leq\dots+\infty$ and $\{\phi_{n}\}$ is an orthonormal basis of $\mathrm{L}^{2}(\Omega)$. Hence $u_{f^{*}}(0,x)\equiv 0$ implies
$e^{-\tilde{T}\lambda_{n}}\langle\phi_{n},f^{*}\rangle=0,\qquad\forall\ n\in\mathbb{N}.$
This further implies $f^{*}\equiv 0$, contradicts $\|f^{*}\|_{\mathrm{L}^{2}(\Omega)}=1$. Hence, for a given $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$.
Second proof: We now provide a quantitative proof based on variational approach.
Consider the functional:
$\mathcal{J}_{\epsilon,a}(f^{*})=\frac{1}{2}\int_{\Omega}|u_{f^{*}}(0,\cdot)|^{2}\,\rm{d}x+\epsilon\|f^{*}\|_{\mathrm{L}^{2}(\Omega)}-\int_{\Omega}hf^{*}\,\rm{d}x,$
where $\epsilon>0$ is a fix quantity and $h\in\mathrm{L}^{2}(\Omega)$ is a given function. Furthermore $u_{f^{*}}$ and $f^{*}$ are connected by (13) with initial data $f^{*}$ at $T=\tilde{T}$. We intend to show that the functional ${\mathcal{J}_{\epsilon,a}:\mathrm{L}^{2}(\Omega)\to\mathbb{R}\cup\{+\infty\}}$ is lower semicontinuous, strictly convex and coercive.
  • (a).
    Continuity: Let $f_{n}^{*}\to f^{*}$ in $\mathrm{L}^{2}(\Omega)$. The solution of (13), $u_{f^{*}_{m}}$ corresponding to the initial data $f_{m}^{*}$ can be expressed as
    $u_{f^{*}_{m}}(t,\cdot)=\sum\limits_{n\in\mathbb{N}}e^{-\lambda_{n}(\tilde{T}-t)}\phi_{n}(x)\langle\phi_{n},f_{m}^{*}\rangle,$
    where $\phi_{n}$ are the eigenfunctions of (6). Hence
    $\|u_{f_{m}^{*}}(0,\cdot)-u_{f^{*}}(0,\cdot)\|^{2}_{\mathrm{L}^{2}(\Omega)}=\sum\limits_{n\in\mathbb{N}}e^{-2\tilde{T}\lambda_{n}}|\langle\phi_{n},f_{m}^{*}\rangle-\langle\phi_{n},f^{*}\rangle|^{2}.$
    From the fact that $f_{n}^{*}\to f^{*}$ in $\mathrm{L}^{2}(\Omega)$, we have that
    ${\sum\limits_{m\in\mathbb{N}}|\langle\phi_{n},f_{m}^{*}\rangle-\langle\phi_{n},f^{*}\rangle|^{2}\to 0}.$
    Hence
    $\|u_{f_{m}^{*}}(0,\cdot)-u_{f^{*}}(0,\cdot)\|^{2}_{\mathrm{L}^{2}(\Omega)}\leq e^{-2\lambda_{1}\tilde{T}}\sum\limits_{n\in\mathbb{N}}|\langle\phi_{n},f_{m}^{*}\rangle-\langle\phi_{n},f*\rangle|^{2}\to 0.$
    This proves continuity of the functional $\mathcal{J}_{\epsilon,a}$.
  • (b).
    Strictly Convexity: It follows from the strict convexity of the the square of the $\mathrm{L}^{2}$-norm.
  • (c).
    Coercivity: Let $f_{m}^{*}\in\mathrm{L}^{2}(\Omega)$, for all $m\in\mathbb{N}$ and ${\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ as $m\to+\infty$. Consider the functions ${g_{m}^{*}=\frac{f_{m}^{*}}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. Then ${\|g_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}=1}$ for all $m\in\mathbb{N}$. Let $u_{g^{*}_{m}}$ be the solution of (13) with respect to the initial data $g_{m}^{*}$. Then ${u_{g^{*}_{m}}=\frac{u_{f^{*}_{m}}}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. We divide the functional $\mathcal{J}_{\epsilon,a}$ by $\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}$. It yields
    $\frac{\mathcal{J}_{\epsilon,a}(f_{m}^{*})}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}=\frac{1}{2}\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\int_{\Omega}|u_{g_{m}^{*}}(0,\cdot)|^{2}\,\rm{d}x+\epsilon-\int_{\Omega}hg_{m}^{*}\,\rm{d}x.$
    There can be two different scenarios here. If ${\liminf\limits_{m\in\mathbb{N}}\int_{\Omega}|u_{g_{m}^{*}}(0,\cdot)|^{2}>0}$, then ${\frac{\mathcal{J}_{\epsilon,a}(f_{m}^{*})}{\|f_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}\to+\infty$ as $m\to+\infty$. The other case when there exists a subsequence (still indexed by $m$) such that ${\liminf\limits_{n\in\mathbb{N}}\int_{\Omega}|u_{g_{m}^{*}}(0,\cdot)|^{2}=0}$. From the compactness of $\mathrm{L}^{2}$ space, we have that $g_{m}^{*}\rightharpoonup g^{*}$ weakly in $\mathrm{L}^{2}(\Omega)$. The functions $u_{g_{m}^{*}}$ can be expressed as
    $u_{g_{m}^{*}}(0,\cdot)=\sum\limits_{n\in\mathbb{N}}e^{-\tilde{T}\lambda_{n}}\phi_{n}\langle\phi_{n},g_{m}^{*}\rangle,$
    where $\phi_{n}$’s are the eigenfunctions corresponding to (6). Let $u_{g^{*}}$ be the solution of (1) with initial data $g^{*}$.
    $u_{g^{*}}(0,\cdot)=\sum\limits_{n\in\mathbb{N}}e^{-\tilde{T}\lambda_{n}}\phi_{n}\langle\phi_{n},g^{*}\rangle.$
    The lower semicontinuity of $\mathrm{L}^{2}$ norm implies ${\|g^{*}\|_{\mathrm{L}^{2}(\Omega)}\leq\|g^{*}_{m}\|_{\mathrm{L}^{2}(\Omega)}\leq 1}$. We will show $u_{g^{*}_{m}}(0,\cdot)\to u_{g^{*}}(0,\cdot)$ converges in $\mathrm{L}^{2}(\Omega)$ strongly.
    $\|u_{g_{m}^{*}}(0,\cdot)-u_{g^{*}}(0,\cdot)\|^{2}_{\mathrm{L}^{2}(\Omega)}=$ $\sum\limits_{n\in\mathbb{N}}e^{-2\tilde{T}\lambda_{n}}|\langle\phi_{n},g^{*}_{m}\rangle-\langle\phi_{n},g^{*}\rangle|^{2}$
    $\leq$ $\sum\limits_{n=1}^{n_{K}}e^{-2\tilde{T}\lambda_{n}}|\langle\phi_{n},g^{*}_{m}\rangle-\langle\phi_{n},g^{*}\rangle|^{2}$
    $+$ $2\sum\limits_{n=n_{K}+1}^{\infty}e^{-2\tilde{T}\lambda_{n}}+2\sum\limits_{n=n_{K}+1}^{\infty}e^{-2\tilde{T}\lambda_{n}}$
    $\leq$ $\sum\limits_{n=1}^{n_{K}}e^{-2\tilde{T}\lambda_{n}}|\langle\phi_{n},g^{*}_{m}\rangle-\langle\phi_{n},g^{*}\rangle|^{2}$
    $+$ $4C_{1}(a)\sum\limits_{n=n_{K}+1}^{\infty}\frac{1}{\lambda_{k}^{\frac{n}{s}}}$
    $\leq$ $\sum\limits_{n=1}^{n_{K}}e^{-2\tilde{T}\lambda_{n}}|\langle\phi_{n},g^{*}_{m}\rangle-\langle\phi_{n},g^{*}\rangle|^{2}$
    $+$ $4C_{1}(a)C_{2}(\Omega,n,a)\sum\limits_{n=n_{K}+1}^{\infty}\frac{1}{k^{2}}.$
    The last inequality follows from one sided Weyl’s law [see [CGCV22], [BG59]]. Here $C_{1}(a)$, $C_{2}(\Omega,n,a)$ are two positive constant depends on the domain, the dimension and the fractional power ‘$a$’. We choose $n_{K}\in\mathbb{N}$ such that ${\sum\limits_{n=n_{K}+1}^{\infty}\frac{1}{k^{2}}\leq\delta}$ for a given $\delta>0$. Hence for a given $\delta>0$, there exists large $m_{K}\in\mathbb{N}$ such that for all $m\geq m_{K}$:
    $\|u_{g_{m}^{*}}(0,\cdot)-u_{g^{*}}(0,\cdot)\|^{2}_{\mathrm{L}^{2}(\Omega)}\leq 2\delta.$
    Hence we deduce ${u_{g_{m}^{*}}(0,\cdot)\to u_{g^{*}}(0,\cdot)}$ strongly in $\mathrm{L}^{2}(\Omega)$, which further implies
    $u_{g^{*}}\equiv 0.$
    Hence ${e^{-\tilde{T}\lambda_{n}}\langle\phi_{n},g^{*}\rangle=0}$ i.e., ${\langle\phi_{n},g^{*}\rangle=0}$ for all $n\in\mathbb{N}$. It yields
    $g^{*}\equiv 0.$
    As $g_{n}^{*}$ converges to $g^{*}$ weakly we have ${\int_{\Omega}hg_{n}^{*}\to 0}$. Hence ${\frac{\mathcal{J}_{\epsilon,a}(f_{n}^{*})}{\|f_{n}^{*}\|_{\mathrm{L}^{2}(\Omega)}}>\frac{\epsilon}{2}}$ for large $m$. Hence we deduce ${\mathcal{J}_{\epsilon,a}(f_{n}^{*})\to+\infty}$ as ${\|f_{n}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$. Hence $\mathcal{J}_{\epsilon,a}$ is coercive.
Continuity, strict convexity and coercivity helps us conclude that $\mathcal{J}_{\epsilon,a}$ attains it’s minimum for some $f_{min}^{*}\in\mathrm{L}^{2}(\Omega)$. We fix a function $g^{*}\in\mathrm{L}^{2}(\Omega)$ and let $u_{g^{*}}$ be the solution of (13) corresponding to the initial data $g^{*}$. We have
$\mathcal{J}_{\epsilon,a}(f_{min}^{*}+\alpha g^{*})-\mathcal{J}_{\epsilon,a}(f_{min}^{*})\geq 0,\quad\forall\,\alpha\in\mathbb{R}.$
For $\alpha\geq 0$, it yields
$\frac{1}{2}\int_{\Omega}2\alpha u_{f^{*}_{min}}(0,\cdot)$ $u_{g^{*}}(0,\cdot)\,\rm{d}x+\frac{1}{2}\int_{\Omega}\alpha^{2}(u_{g^{*}}(0,\cdot))^{2}\,\rm{d}x$
$+$ $\epsilon\frac{\alpha^{2}\|g^{*}\|^{2}_{\mathrm{L}^{2}(\Omega)}+2\alpha\int_{\Omega}f^{*}_{min}g^{*}\,\rm{d}x}{\|f^{*}+\alpha g^{*}\|_{\mathrm{L}^{2}(\Omega)}+\|f_{min}^{*}\|_{\mathrm{L}^{2}(\Omega)}}-\alpha\int_{\Omega}hg^{*}\,\rm{d}x\geq 0.$
We divide the whole expression by $\alpha$ and let $\alpha\to 0\uparrow$. It yields
$\int_{\Omega}u_{f^{*}_{min}}(0,\cdot)u_{g^{*}}(0,\cdot)\,\rm{d}x+\epsilon\frac{\int_{\Omega}f^{*}_{min}g^{*}\,\rm{d}x}{\|f_{min}^{*}\|_{\mathrm{L}^{2}(\Omega)}}-\int_{\Omega}hg^{*}\,\rm{d}x\geq 0.$
Employing Cauchy-Schwartz inequality, we obtain
(19) $\int_{\Omega}u_{f^{*}_{min}}(0,\cdot)u_{g^{*}}(0,\cdot)\,\rm{d}x-\int_{\Omega}hg^{*}\,\rm{d}x\geq-\epsilon\|g^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
Similarly for $\alpha\leq 0$, we obtain
(20) $\int_{\Omega}u_{f^{*}_{min}}(0,\cdot)u_{g^{*}}(0,\cdot)\,\rm{d}x-\int_{\Omega}hg^{*}\,\rm{d}x\leq\epsilon\|g^{*}(0,\cdot)\|_{\mathrm{L}^{2}(\Omega)}.$
Together, we obtain
(21) $\left|\int_{\Omega}u_{f^{*}_{min}}(0,\cdot)u_{g^{*}}(0,\cdot)\,\rm{d}x-\int_{\Omega}hg^{*}\,\rm{d}x\right|\leq\epsilon\|g^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
We choose $f=\eta u_{f^{*}_{min}}(0,\cdot)$ for the initial condition in (1), where $\eta\in C_{c}^{\infty}(\Omega)$ to be choosen later. Let $f_{1}=(1-\eta)u_{f^{*}_{min}}(0,\cdot)$. Let $u_{f},u_{f_{1}}$ be the solutions to (1) with initial datum $f$ and $f_{1}$ respectively. Thanks to (9), we can conclude $f\in C_{c}^{\infty}(\Omega)$. [see [FRRO16], [Gru23]]. Integration by parts (17) yields that:
$\int_{\Omega}fu_{g^{*}}(0,\cdot)\,\rm{d}x$ $=\int_{\Omega}u_{f}(\tilde{T},\cdot)g^{*}\,\rm{d}x$
$\int_{\Omega}f_{1}u_{g^{*}}(0,\cdot)\,\rm{d}x$ $=\int_{\Omega}u_{f_{1}}(\tilde{T},\cdot)g^{*}\,\rm{d}x.$
Substituting this in (21), we obtain
$\left|\int_{\Omega}\big{(}u_{f}(\tilde{T},\cdot)+u_{f_{1}}(\tilde{T},\cdot)-h\big{)}g^{*}\,\rm{d}x\right|\leq\epsilon\|g^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
Since $g^{*}\in\mathrm{L}^{2}(\Omega)$ is arbitrary, we obtain:
$\|u_{f}(\tilde{T},\cdot)+u_{f_{1}}(\tilde{T},\cdot)-h\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon.$
Triangle inequality yields:
(22) $\|u_{f}(\tilde{T},\cdot)-h\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon+\|u_{f_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}.$
It remains to choose $\eta\in C_{c}^{\infty}(\Omega)$ such that ${\|u_{f_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon}$. Thanks to (7), we have
$\sup\limits_{t\geq T}\|u_{f_{1}}(\tilde{T},\cdot)\|_{C_{c}^{a}(\mathbb{R}^{N})}\leq C_{1}(\tilde{T})\|f_{1}\|_{\mathrm{L}^{2}(\Omega)}\leq C_{1}(\tilde{T})\|(1-\eta)u_{f^{*}_{min}}(0,\cdot)\|_{\mathrm{L}^{2}(\Omega)}.$
Again, thanks to (7), we have
(23) $\|u_{f^{*}_{min}}(0,\cdot)\|_{C^{a}_{c}(\mathbb{R}^{N})}\leq C^{*}_{1}(\tilde{T})\|f^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}.$
If $f^{*}_{min}\equiv 0$, then thanks to (23), we can choose arbitrary $\eta$. So Let $f^{*}_{min}\neq 0$, and $\tilde{\Omega}$ is a domain compactly contained in $\Omega$ such that
$\mbox{Volume}\left|\Omega\setminus\tilde{\Omega}\right|\leq\epsilon^{2}C_{1}^{-2}(\tilde{T})\left(C_{1}^{*}\right)^{-2}(\tilde{T})\|f^{*}_{min}\|^{-2}_{\mathrm{L}^{2}(\Omega)}.$
Employing Hölder inequality, we obtain
$\|u_{f_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq C_{1}(\tilde{T})\sup\{|1-\eta|\}\sup\|u_{f_{min}^{*}}(0,\cdot)\|_{C^{0}_{c}(\mathbb{R}^{N})}\left(\mbox{Volume}\left|\Omega\setminus\tilde{\Omega}\right|\right)^{\frac{1}{2}},$
where $\eta\in C_{c}^{\infty}(\Omega)$ such that $\eta\equiv 1$ in $\tilde{\Omega}$. The choice of $\tilde{\Omega}$ yields
$\|u_{f_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon.$
Hence for a given $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$. ∎
Next we are interested in singular boundary data system (2). We like to analyze the density in $\mathrm{L}^{2}(\Omega)$ space on a particular time slice by varying singular boundary data. We start with an observability estimate.
Proposition 9.
Let $f^{*}\in L^{2}(\Omega)$ and let $0\leq t_{1}<t_{2}\leq T$. Let $u_{f^{*}}$ be the solution of (13) with initial data $f^{*}$. Then the following observability estimate holds
$\int_{\Omega}u_{f^{*}}^{2}(t_{1},\cdot)\,\rm{d}x-$ $\int_{\Omega}u_{f^{*}}^{2}(t_{2},\cdot)\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x\,\rm{d}t$
$=$ $-\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}t.$
Furthermore there exists a positive constant $\mathcal{A}$, depends only on the domain, the dimension and the exponent ‘$a$’, such that:
$\mathcal{A}\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}s\geq\|u_{f^{*}}\|_{\mathrm{L}^{2}((t_{1},t_{2})\times\Omega)}.$
Proof.
Thanks to (9), we have $\nabla u_{f^{*}},(-\Delta)^{a}u_{f^{*}}\in\mathrm{L}^{2}(\Omega)$ for all $t\in(0,T)$ and thanks to (7),(8),(10),(11) we can apply fractional Pohozaev identity (12) for the function $u_{f^{*}}$. It yields
$\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $u_{f^{*}}(-\Delta)^{a}u_{f^{*}}\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
Equation (13) yields
(24) $\int_{\Omega}(x\cdot\nabla u_{f^{*}})$ $(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x$
-- $\frac{2a-n}{2}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
We intend to calculate the term ${\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x}$. Gauss divergence theorem yields:
$\int_{\Omega}(x\cdot\nabla u_{f^{*}})$ $(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=-\int_{\Omega}\nabla\cdot\left(x(-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x$
$=$ $-\int_{\Omega}n(-\Delta)^{a}u_{f^{*}}u_{f^{*}}\,\rm{d}x-\int_{\Omega}x\cdot\nabla\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x.$
Using equation (13), we obtain
(25) $\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=$ $-n\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x+n\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x$
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x.$
We compute the last term of the right hand side above:
$-\int_{\mathbb{R}^{N}}x\cdot\nabla$ $\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x=-\int_{\mathbb{R}^{N}}\nabla_{\xi}\cdot\left(\xi|\xi|^{2a}\hat{u_{f^{*}}}\right)\hat{u_{f^{*}}}\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}\left(n|\xi|^{2a}\hat{u_{f^{*}}}+2a|\xi|^{2a}\hat{u_{f^{*}}}+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\hat{u_{f^{*}}}\right)\right)\hat{u_{f^{*}}}\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}n|\xi|^{2a}\hat{u_{f^{*}}}^{2}+2a|\xi|^{2a}\hat{u_{f^{*}}}^{2}+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\frac{\hat{u_{f^{*}}}^{2}}{2}\right)\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}(n+2a)|\xi|^{2a}\hat{u_{f^{*}}}^{2}\,\rm{d}\xi+\int_{\mathbb{R}^{N}}\nabla_{\xi}.\left(\xi|\xi|^{2a}\right)\frac{\hat{u_{f^{*}}}^{2}}{2}\,\rm{d}\xi$
$=$ $-\frac{n+2a}{2}\int_{\mathbb{R}^{N}}|\xi|^{2a}\hat{u_{f^{*}}}^{2}\,\rm{d}\xi=-\frac{n+2a}{2}\int_{\Omega}(-\Delta)^{a}u_{f^{*}}u_{f^{*}}\,\rm{d}x$
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}u_{f^{*}}\right)u_{f^{*}}\,\rm{d}x=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}u_{f^{*}}u_{f^{*}}-qu_{f^{*}}^{2}\right)\,\rm{d}x$
$=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)-qu_{f^{*}}^{2}\right)\,\rm{d}x.$
Substituting this in (25), we obtain
$\int_{\Omega}(x\cdot\nabla u_{f^{*}})(-\Delta)^{a}u_{f^{*}}\,\rm{d}x=-\frac{3n+2a}{2}\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x+\frac{3n+2a}{2}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x.$
Substituting this in (24) yields
$(n+2a)\int_{\Omega}\partial_{t}\left(\frac{u_{f^{*}}^{2}}{2}\right)\,\rm{d}x-(n+2a)\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma=0.$
Integrating with respect to time yields:
(26) $\int_{\Omega}u_{f^{*}}^{2}(t_{1},\cdot)\,\rm{d}x-$ $\int_{\Omega}u_{f^{*}}^{2}(t_{2},\cdot)\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}qu_{f^{*}}^{2}\,\rm{d}x\,\rm{d}t$
$=$ $-\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}t.$
We recall equation (13):
(27) $-\partial_{t}u_{f^{*}}+(-\Delta)^{a}u_{f^{*}}+qu_{f^{*}}=0$
with $u_{f^{*}}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$. We multiply the equation (27) by $u_{f^{*}}$ and integrate with respect to space time.
$\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{2},\cdot)\,\rm{d}x+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}u_{f*}$ $(-\Delta)^{a}u_{f^{*}}\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}qu^{2}_{f*}\,\rm{d}x\,\rm{d}t=0.$
This implies
$\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{2},\cdot)\,\rm{d}x+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}$ $\left|(-\Delta)^{\frac{a}{2}}u_{f^{*}}\right|^{2}\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}qu^{2}_{f*}\,\rm{d}x\,\rm{d}t=0,$
i.e.
$\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{u_{f^{*}}^{2}}{2}(t_{2},\cdot)\,\rm{d}x+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}qu^{2}_{f*}\,\rm{d}x\,\rm{d}t$
$=-\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}u_{f^{*}}\right|^{2}\,\rm{d}x\,\rm{d}t.$
By Hardy-Littlewood-Sobolev estimate:
$\|u_{f^{*}}\|_{\mathrm{L}^{2}(\Omega)}\leq C_{Vol}\|u_{f^{*}}\|_{\mathrm{L}^{\frac{2n}{n-a}}(\Omega)}\leq C_{n,a,\Omega}\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}u_{f^{*}}\right|^{2}\,\rm{d}x\,\rm{d}t,$
where $C_{Vol}$ and $C_{n,a,\Omega}$ are two fixed positive constants depends on the domain and the dimension only. Let ${\mathcal{A}=\frac{1}{2C_{n,a,\Omega}}}$. Then from (26), we conclude
$\mathcal{A}\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{u_{f^{*}}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}s\geq\|u_{f^{*}}\|_{\mathrm{L}^{2}((t_{1},t_{2})\times\Omega)}.$
We now move onto the proof of our next density result.
Proof of theorem 3:
Proof.
Again we present two proofs here. One using integration by parts (14) and another by a variational approach.
First proof: Let the set ${\{u_{F}(\tilde{T},\cdot),\,F\in C_{c}^{\infty}((0,T)\times\Omega)\}}$ is not dense in $\mathrm{L}^{2}(\Omega)$. Then by Hahn-Banach theorem there exists $g\in\mathrm{L}^{2}(\Omega)$ such that $\|g\|_{L^{2}(\Omega)}=1$ and
(28) $\langle u_{F}(\tilde{T},\cdot),g\rangle=0,\qquad\forall F\in C_{c}^{\infty}((0,T)\times\Omega).$
Let $u_{g}\in\mathrm{L^{2}((0,\tilde{T});H^{a}(\Omega))}$ be the solution of the adjoint system (13) with $g$ as initial data at the time $T=\tilde{T}$. Theorem 8 ensures ${u_{F}\in\mathrm{L}^{2}\left((0,\tilde{T});H^{(a-1)(2a)}(\overline{\Omega})\right)}$ respectively. We apply integration by parts formula (16).
$\int_{\Omega}u_{g}(-\Delta)^{a}u_{F}-\int_{\Omega}u_{F}(-\Delta)^{a}u_{g}=-\Gamma(s)\Gamma(s-1)\int_{\partial\Omega}\frac{u_{F}}{d^{s-1}}\frac{u_{g}}{d^{s}}.$
From equation (2), we obtain
$\int_{\Omega}\partial_{t}u_{g}u_{F}+u_{g}\partial_{t}u_{F}=\int_{\Omega}\partial_{t}\left(u_{g}u_{F}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{s}}.$
Integrating with respect to $t$ yields:
$\int_{\Omega}u_{g}(\tilde{T},\cdot)u_{F}(\tilde{T},\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{a}}\,\rm{d}x\,\rm{d}s.$
Thanks to (28) and since $F\in C_{c}^{\infty}((0,T)\times\Omega)$ is arbitrary, we deduce that
$\frac{u_{g}}{d^{a}}\equiv 0.$
Observability estimate (9) yields
$\|u_{g}\|_{\mathrm{L}^{2}((0,\tilde{T})\times\Omega)}=0\implies u_{g}\equiv 0\implies g\equiv 0.$
Next we present a proof via variational approach. It’s more quantitative in nature.
Second proof: Consider the functional
$\mathcal{J}_{\epsilon,a}(F^{*})=\frac{1}{2}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta^{2}_{\tilde{T}}\left|\frac{u_{F^{*}}}{d^{a}}\right|^{2}\,\rm{d}x\,\rm{d}t+\epsilon\|{F^{*}}\|_{\mathrm{L}^{2}(\Omega)}-\int_{\Omega}h{F^{*}}\,\rm{d}x,$
where $\epsilon>0$ is a fix quantity, $h\in\mathrm{L}^{2}(\Omega)$ is a given function and $\eta_{\tilde{T}}\in C_{c}^{\infty}((0,\tilde{T}))$ such that $\eta_{\tilde{T}}\equiv 1$ in $[T_{1},T_{2}]\subset(0,\tilde{T})$. Furthermore $u_{F^{*}}$ and $F^{*}$ are connected by (13) with $F^{*}$ as the initial data at $T=\tilde{T}$. We intend to show the functional ${\mathcal{J}_{\epsilon,a}:\mathrm{L}^{2}(\Omega)\to\mathbb{R}\cup\{+\infty\}}$ is lower semicontinuous, strictly convex and coercive. Just like the case of theorem 2 continuity of $\mathcal{J}_{\epsilon,a}$ follows from the eigenvalue representation and (8). Here (8) provides the boundary continuity. Strict convexity follows from the strict convexity of the $\mathrm{L}^{2}$-norm and observability estimate (theorem 9). We only prove coercivity of the functional $\mathcal{J}_{\epsilon,a}$.
  • (a).
    Coercivity: Let $F_{m}^{*}\in\mathrm{L}^{2}(\Omega)$ for all $m\in\mathbb{N}$ and ${\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ as $m\to+\infty$. Consider the functions ${G_{m}^{*}=\frac{F_{m}^{*}}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. Then ${\|G_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}=1}$ for all $m\in\mathbb{N}$. Let $u_{G^{*}_{m}}$ be the solution of (13) with respect to the initial data $G_{m}^{*}$. Then ${u_{G^{*}_{m}}=\frac{u_{F^{*}_{m}}}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}}$. We divide the functional $\mathcal{J}_{\epsilon,a}$ by $\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}$. It yields
    $\frac{\mathcal{J}_{\epsilon,a}(F^{*}_{m})}{\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}}=\frac{1}{2}\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\left|\frac{u_{G_{m}^{*}}}{d^{a}}\right|^{2}\,\rm{d}x\,\rm{d}t+\epsilon-\int_{\Omega}h{G_{m}^{*}}\,\rm{d}x.$
    There can be two different scenarios. One is
    $\liminf_{m\in\mathbb{N}}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\left|\frac{u_{G_{m}^{*}}}{d^{a}}\right|^{2}\,\rm{d}x\,\rm{d}t>0\implies\mathcal{J}_{\epsilon,a}\to+\infty.$
    The second scenario is ${\liminf_{m\in\mathbb{N}}\int_{0}^{\tilde{T}}\int_{\partial\Omega}\eta_{\tilde{T}}^{2}\left|\frac{u_{G_{m}^{*}}}{d^{s}}\right|^{2}\,\rm{d}x\,\rm{d}t=0}$. Thanks to the observability estimate (9) we have that ${\|u_{G_{m}^{*}}\|_{\mathrm{L}^{2}((T_{1},T_{2})\times\Omega)}\to 0}$ as $m\to+\infty$ for all $T_{1}<T_{2}\in[0,\tilde{T}]$. We represents the solution $u_{G_{m}^{*}}$ as
    $u_{G_{m}^{*}}(t,x)=\sum\limits_{n\in\mathbb{N}}e^{-\lambda_{n}(\tilde{T}-t)}\langle\phi_{n},G_{m}^{*}\rangle\phi_{n}(x).$
    Hence ${\langle\phi_{n},G_{m}^{*}\rangle\to 0}$ as $m\to+\infty$. As $G_{m}^{*}$ is uniformly bounded in $\mathrm{L}^{2}$, there exists $G^{*}\in\mathrm{L}^{2}(\Omega)$ such $G_{m}^{*}\rightharpoonup G^{*}$ weakly in $\mathrm{L}^{2}(\Omega)$. This further implies
    $\langle\phi_{n},G_{m}^{*}\rangle\to\langle\phi_{n},G^{*}\rangle\to 0\ \ \mbox{as}\ m\to+\infty.$
    Hence $G^{*}\equiv 0$ and we have that
    $\int_{\Omega}hG_{m}^{*}\to 0\ \ \mbox{as}\ m\to+\infty.$
    Hence $\mathcal{J}_{\epsilon,a}(F_{m}^{*})\to+\infty$ as ${\|F_{m}^{*}\|_{\mathrm{L}^{2}(\Omega)}\to+\infty}$ i.e., $\mathcal{J}_{\epsilon,a}$ is coercive.
Continuity, strict convexity and coercivity helps us conclude that $\mathcal{J}_{\epsilon,a}$ attains it’s minimum for some $F_{min}^{*}\in\mathrm{L}^{2}(\Omega)$. We fix a function $G^{*}\in\mathrm{L}^{2}(\Omega)$ and let $u_{G^{*}}$ be the solution of (13) corresponding to the initial data $G^{*}$. We have
$\mathcal{J}_{\epsilon,a}(F_{min}^{*}+\alpha G^{*})-\mathcal{J}_{\epsilon,a}(F_{min}^{*})\geq 0,\quad\forall\,\alpha\in\mathbb{R}.$
For $\alpha\geq 0$, it yields
$\frac{1}{2}$ $\int_{0}^{\tilde{T}}\int_{\Omega}2\alpha\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t+\frac{1}{2}\int_{0}^{\tilde{T}}\int_{\Omega}\alpha^{2}\eta_{\tilde{T}}^{2}\frac{(u_{G^{*}})^{2}}{d^{2a}}\,\rm{d}x\,\rm{d}t$
$+$ $\epsilon\frac{\alpha^{2}\|G^{*}\|^{2}_{\mathrm{L}^{2}(\Omega)}+2\alpha\int_{\Omega}F^{*}_{min}{G^{*}}\,\rm{d}x}{\|F^{*}_{min}+\alpha G^{*}\|_{\mathrm{L}^{2}(\Omega)}+\|F_{min}^{*}\|_{\mathrm{L}^{2}(\Omega)}}-\alpha\int_{\Omega}hG^{*}\,\rm{d}x\geq 0.$
We divide the whole expression by $\alpha$ and let $\alpha\to 0\uparrow$. It yields
$\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t+\epsilon\frac{\int_{\Omega}F^{*}_{min}G^{*}\,\rm{d}x}{\|F_{min}^{*}\|_{\mathrm{L}^{2}(\Omega)}}-\int_{\Omega}hG^{*}\,\rm{d}x\geq 0.$
Hence employingl Cauchy-Schwartz inequality, we obtain
$\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\geq-\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
Similarly for $\alpha\leq 0$, we obtain
$\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
Together, we obtain:
(29) $\left|\int_{0}^{\tilde{T}}\int_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}t-\int_{\Omega}hG^{*}\,\rm{d}x\right|\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
We choose ${F=\eta_{\Omega}\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}}$ for the singular boundary data in (2), where $\eta_{\Omega}\in C_{c}^{\infty}(\partial\Omega)$ to be chosen later. Let ${F_{1}=(1-\eta_{\Omega})\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}}$. Let $u_{F},u_{F_{1}}$ be the solutions to (2) with singular boundary datum $F$ and $F_{1}$ respectively. Thanks to (9), we can conclude $F\in C_{c}^{\infty}((0,T)\times\Omega)$. [see [FRRO16], [Gru23]]. Integration by parts (16) yields that:
$\int_{\Omega}u_{F}(\tilde{T},\cdot)G^{*}(\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s.$
$\int_{\Omega}u_{F_{1}}(\tilde{T},\cdot)G^{*}(\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}\frac{u_{F_{1}}}{d^{a-1}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s.$
Substituting this in (29), we obtain
$\left|\int_{\Omega}\big{(}u_{F}(\tilde{T},\cdot)+u_{F_{1}}(\tilde{T},\cdot)-h\big{)}G^{*}\,\rm{d}x\right|\leq\epsilon\|G^{*}\|_{\mathrm{L}^{2}(\Omega)}.$
Since $G^{*}\in\mathrm{L}^{2}(\Omega)$ is arbitrary, we obtain:
$\|u_{F}(\tilde{T},\cdot)+u_{F_{1}}(\tilde{T},\cdot)-h\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon.$
[…]
$\frac{u_{F^{*}_{min}}}{d^{s}}(t,z)=\sum\limits_{n\in\mathbb{N}}e^{-(\tilde{T}-t)\lambda_{n}}\langle\phi_{n},F^{*}_{min}\rangle\frac{\phi_{n}}{d^{s}}(z),\quad\forall z\in\partial\Omega.$
We use the fact that ${\left|\frac{\phi_{n}}{d^{s}}\right|\leq C_{\Omega}\lambda_{n}^{w}}$, where $w\in\mathbb{N}$ and $C_{\Omega}$ two positive constants depends on the domain, the dimension and the exponent ‘$a$’ only [see [AGCV23], [CGCV21], [Gru15b], [CGCV22], [FRRO16], [CKS10]]. It yields
$\left\|\frac{u_{F^{*}_{min}}}{d^{a}}\right\|_{\mathrm{L}^{\infty}([T_{3},T_{4}]\times\partial\Omega)}\leq$ $C_{\Omega}\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}\sum\limits_{n\in\mathbb{N}}e^{-\lambda_{n}(\tilde{T}-T_{4})}\lambda_{n}^{w}$
$\leq$ $C_{\Omega}\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}C_{\tilde{T},T_{4},w},$
where $C_{\tilde{T},T_{4},w}$ a positive constant depends only on $\tilde{T}-T_{4}$ and $w$. Let $u_{G^{*}}$ be the solution of (13) with initial data $G^{*}$. Integration by parts (16) yields:
$\int_{\Omega}u_{F_{1}}(\tilde{T},\cdot)G^{*}\,\rm{d}x=$ $-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}\frac{u_{F_{1}}}{d^{a-1}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s$
$=$ $-\Gamma(a)\Gamma(a-1)\int_{0}^{\tilde{T}}\int_{\partial\Omega}(1-\eta_{\Omega})\eta_{\tilde{T}}^{2}\frac{u_{F^{*}_{min}}}{d^{a}}\frac{u_{G^{*}}}{d^{a}}\,\rm{d}x\,\rm{d}s.$
Employing Hölder inequality, we obtain
$\int_{\Omega}u_{F_{1}}(\tilde{T},\cdot)G^{*}(\cdot)\,\rm{d}x\leq$ $\mathcal{Z}\text{Vol}(|\text{supp}(1-\eta_{\Omega})|)\left\|\frac{u_{F^{*}_{min}}}{d^{a}}\right\|_{\mathrm{L}^{\infty}([T_{3},T_{4}]\times\partial\Omega)}$
$\times$ $\left\|\frac{u_{G^{*}_{min}}}{d^{a}}\right\|_{\mathrm{L}^{\infty}([T_{3},T_{4}]\times\partial\Omega)}$
$\leq$ $\mathcal{Z}\text{Vol}(|\text{supp}(1-\eta_{\Omega})|)\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}\|G^{*}\|_{\mathrm{L}^{2}(\Omega)},$
where ${\mathcal{Z}:=\tilde{T}\big{|}\Gamma(s)\Gamma(s-1)\big{|}C_{\Omega}^{2}C_{\tilde{T},T_{4},w}^{2}}$. Since $G^{*}$ is arbitrary we have that
$\|u_{F_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq\mathcal{Z}\text{Vol}(\text{supp}|1-\eta_{\Omega}|)\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}.$
If $F_{min}^{*}\equiv 0$, then ${\|u_{F_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}=0}$. Hence we are done. Otherwise, let $\eta_{\Omega}\in C_{c}^{\infty}(\partial\Omega)$ such that $\eta\equiv 1$ in $\tilde{\partial\Omega}$ compactly contained in $\partial\Omega$ such that
$\text{Vol}\left(\left|\partial\Omega\setminus\tilde{\partial\Omega}\right|\right)\leq\frac{\epsilon}{\mathcal{Z}\|F^{*}_{min}\|_{\mathrm{L}^{2}(\Omega)}}.$
Hence, ${\|u_{F_{1}}(\tilde{T},\cdot)\|_{\mathrm{L}^{2}(\Omega)}\leq\epsilon}$. Thus we conclude our proof. ∎
We next move onto the proofs of the Calderón type inverse problems for the non-local heat equations. We recover the potentials by analyzing the solution at a particular time slice for (1) and by analyzing the solution on the boundary for (2).
3. Calderón type Inverse Problems:
Proof of theorem 4:
Proof.
We show that proving theorem 4 is same as proving theorem 5. Let $u_{1}$, $u_{2}$ be the solution of (4). Let for $i=1,2$, $w_{i}$ satisfies
$-\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{w_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ w_{i}(T,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $F\in C_{c}^{\infty}((0,T)\times\Omega)$. Since ${w_{i}\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)}$ for $i=1,2$ [see theorem 8, theorem 7], integration by parts (16) yields
$\int_{\Omega}u_{i}(-\Delta)^{a}w_{i}-\int_{\Omega}w_{i}(-\Delta)^{a}u_{i}=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{w_{i}}{d^{a-1}}\frac{u_{i}}{d^{a}},\quad i=1,2,$
where, for $i=1,2$, $u_{i}$’s are the solutions of (4). Replacing $(-\Delta)^{a}u_{i}$ and $(-\Delta)^{a}w_{i}$ using the equations yields:
$\int_{\Omega}\partial_{t}u_{i}w_{i}+w_{i}\partial_{t}u_{i}=\int_{\Omega}\partial_{t}\left(u_{i}w_{i}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{w_{i}}{d^{a-1}}\frac{u_{i}}{d^{a}},\quad i=1,2.$
Integrating with respect to $t$ yields:
$\int_{\Omega}w_{i}(0,\cdot)u_{i}(0,\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{T}\int_{\partial\Omega}\frac{w_{i}}{d^{a-1}}\frac{u_{i}}{d^{a}}\,\rm{d}x\,\rm{d}s,\quad i=1,2.$
Subtracting we obtain
$\int_{\Omega}f(w_{1}(0,\cdot)-w_{2}(0,\cdot))\,\rm{d}x=0,$
as we have ${\frac{u_{1}}{d^{a}}=\frac{u_{2}}{d^{a}}}$ for all $f\in C_{c}^{\infty}(\Omega)$. Hence we obtain
$w_{1}(0,\cdot)\equiv w_{2}(0,\cdot).$
Our goal is to obtain $q_{1}\equiv q_{2}$. So our problem boils down to the following problem: Let $w_{i}$, for $i=1,2$ satisfy
$-\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\overline{\Omega}\\ \lim\limits_{\begin{subarray}{c}x\to y\\ y\in\partial\Omega\end{subarray}}\frac{w_{i}}{d^{a-1}}(y)=&F\qquad\qquad\mbox{on}\ (0,T)\times\partial\Omega\\ w_{i}(T,\cdot)=&0\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $F\in C_{c}^{\infty}((0,T)\times\Omega)$. Let
$w_{1}(0,\cdot)\equiv w_{2}(0,\cdot),\quad\forall\,F\in C_{c}^{\infty}((0,T)\times\Omega).$
Then
$q_{1}\equiv q_{2}.$
The change of variable $t\to T-t$ transforms the problem into the one described in Theorem (5). Hence it is enough to prove theorem (5). ∎
Proof of theorem (5):
Proof.
Let For $i=1,2$, let $w_{i}$ satisfies:
$-\partial_{t}w_{i}+(-\Delta)^{a}w_{i}+q_{i}w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ w_{i}=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ w_{i}(T,\cdot)=&f\qquad\qquad\mbox{in}\ \Omega,\end{aligned}\right.$
where $f\in C_{c}^{\infty}(\Omega)$. Since ${u_{i}\in\mathrm{L}^{2}\left((0,T);H^{(a-1)(2a)}(\overline{\Omega})\right)}$ for $i=1,2$ [see theorem 8, theorem 7], integration by parts (16) yields
$\int_{\Omega}w_{i}(-\Delta)^{a}u_{i}-\int_{\Omega}u_{i}(-\Delta)^{a}w_{i}=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{i}}{d^{a-1}}\frac{w_{i}}{d^{a}},\quad i=1,2,$
where, for $i=1,2$, $u_{i}$’s are the solutions of (5). Replacing $(-\Delta)^{a}u_{i}$ and $(-\Delta)^{a}w_{i}$ using the equations yields:
$\int_{\Omega}\partial_{t}u_{i}w_{i}+w_{i}\partial_{t}u_{i}=\int_{\Omega}\partial_{t}\left(u_{i}w_{i}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{i}}{d^{a-1}}\frac{w_{i}}{d^{a}},\quad i=1,2.$
Integrating with respect to $t$ yields:
$\int_{\Omega}w_{i}(T,\cdot)u_{i}(T,\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{T}\int_{\partial\Omega}\frac{u_{i}}{d^{a-1}}\frac{w_{i}}{d^{a}}\,\rm{d}\sigma\,\rm{d}t,\quad i=1,2,$
Substracting, we obtain
$\int_{\Omega}w_{1}(T,\cdot)u_{1}(T,\cdot)-w_{2}(T,\cdot)u_{2}(T,\cdot)\,\rm{d}x=-\Gamma(a)\Gamma(a-1)\int_{0}^{T}\int_{\partial\Omega}F\left(\frac{w_{1}}{d^{a}}-\frac{w_{2}}{d^{a}}\right)\,\rm{d}\sigma\,\rm{d}t.$
As $u_{1}(T,\cdot)\equiv u_{2}(T,\cdot)$, the above relation implies
(30) $\int_{0}^{T}\int_{\partial\Omega}F\left(\frac{w_{1}}{d^{a}}-\frac{w_{2}}{d^{a}}\right)\,\rm{d}\sigma\,\rm{d}t=0\implies\frac{w_{1}}{d^{a}}=\frac{w_{2}}{d^{a}}:=G(\text{say})$
as $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ arbitrary. The difference $w_{1}-w_{2}$ satisfies
$-\partial_{t}(w_{1}-w_{2})+(-\Delta)^{a}(w_{1}-w_{2})+&(q_{1}w_{1}-q_{2}w_{2})=0\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ (w_{1}-w_{2})=&0\qquad\qquad\ \ \quad\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{n}\setminus\Omega\\ (w_{1}-w_{2})(T,\cdot)=&0\qquad\qquad\qquad\qquad\quad\ \ \qquad\mbox{in}\ \Omega,\end{aligned}\right.$
Thanks to (8), (10), (11), we can apply Pohozaev identity (12). It yields
$\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2})(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $(w_{1}-w_{2})(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{w_{1}-w_{2}}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
Equation corresponding to $w_{1}-w_{2}$ yields:
(31) $\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))$ $(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x$
-- $\frac{2a-n}{2}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{(w_{1}-w_{2})}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma.$
We intend to calculate the term ${\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x}$. Gauss divergence theorem yields:
$\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))$ $(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=-\int_{\Omega}\nabla\cdot\left(x(-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x$
$=$ $-\int_{\Omega}n(-\Delta)^{a}(w_{1}-w_{2})(w_{1}-w_{2})\,\rm{d}x-\int_{\Omega}x\cdot\nabla\left((-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x.$
Again equation corresponding to $w_{1}-w_{2}$ yields
(32) $\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=$ $-n\int_{\Omega}\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x+n\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x$
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x.$
We compute the last term of the right hand side above:
$-\int_{\mathbb{R}^{N}}x\cdot\nabla$ $\left((-\Delta)^{a}(w_{1}-w_{2})\right)(w_{1}-w_{2})\,\rm{d}x=-\int_{\mathbb{R}^{N}}\nabla_{\xi}\cdot\left(\xi|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})\right)(\hat{w_{1}}-\hat{w_{2}})\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}\left(n|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})+2a|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\hat{w_{1}}-\hat{w_{2}}\right)\right)(\hat{w_{1}}-\hat{w_{2}})\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}n|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}+2a|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}+\xi|\xi|^{2a}\nabla_{\xi}\cdot\left(\frac{(\hat{w_{1}}-\hat{w_{2}})^{2}}{2}\right)\,\rm{d}\xi$
$=$ $-\int_{\mathbb{R}^{N}}(n+2a)|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}\,\rm{d}\xi+\int_{\mathbb{R}^{N}}\nabla_{\xi}.\left(\xi|\xi|^{2a}\right)\frac{(\hat{w_{1}}-\hat{w_{2}})^{2}}{2}\,\rm{d}\xi$
$=$ $-\frac{n+2a}{2}\int_{\mathbb{R}^{N}}|\xi|^{2a}(\hat{w_{1}}-\hat{w_{2}})^{2}\,\rm{d}\xi=-\frac{n+2a}{2}\int_{\Omega}(-\Delta)^{a}(w_{1}-w_{2})(w_{1}-w_{2})\,\rm{d}x$
Replacing $(-\Delta)^{a}(w_{1}-w_{2})$ in terms of time derivative yields
$-\int_{\mathbb{R}^{N}}x\cdot\nabla\left((-\Delta)^{a}(w_{1}-w_{2})\right)$ $(w_{1}-w_{2})\,\rm{d}x$
$=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}(w_{1}-w_{2})(w_{1}-w_{2})-(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\right)\,\rm{d}x$
$=$ $-\frac{n+2a}{2}\int_{\Omega}\left(\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)-(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\right)\,\rm{d}x.$
Substituting this in (32), we obtain
$\int_{\Omega}(x\cdot\nabla(w_{1}-w_{2}))(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x=-\frac{3n+2a}{2}\int_{\Omega}$ $\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x$
$+\frac{3n+2a}{2}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x.$
Substituting this in (31) yields
$(n+2a)\int_{\Omega}\partial_{t}\left(\frac{(w_{1}-w_{2})^{2}}{2}\right)\,\rm{d}x-(n+2a)\int_{\Omega}$ $(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{(w_{1}-w_{2})}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma=0.$
Integrating with respect to time yields:
$\int_{\Omega}(w_{1}-w_{2})^{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}(w_{1}-w_{2})^{2}(t_{2},\cdot)$ $\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})\,\rm{d}x\rm{d}t$
$=$ $-\Gamma(1+a)^{2}\frac{1}{n+2a}\int_{t_{1}}^{t_{2}}\int_{\partial\Omega}\left(\frac{(w_{1}-w_{2})}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma\,\rm{d}t.$
The equality (30) yields:
(33) $\int_{\Omega}(w_{1}-w_{2})^{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}(w_{1}-w_{2})^{2}(t_{2},\cdot)$ $\,\rm{d}x+2\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})\,\rm{d}x\,\rm{d}t=0$
We recall equation corresponding to $w_{1}-w_{2}$:
$-\partial_{t}(w_{1}-w_{2})+(-\Delta)^{a}(w_{1}-w_{2})+(q_{1}w_{1}-q_{2}w_{2})=0$
with $w_{1}-w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$. We multiply the equation corresponding to $(w_{1}-w_{2})$ by $(w_{1}-w_{2})$ and integrate with respect to space and time.
$\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{2},\cdot)\,\rm{d}x$ $+\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}(w_{1}-w_{2})(-\Delta)^{a}(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t=0.$
This implies
$\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{2},\cdot)\,\rm{d}x+$ $\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}(w_{1}-w_{2})\right|^{2}\,\rm{d}x\,\rm{d}t$
$+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t=0,$
Thanks to (33), we obtain
$\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{1},\cdot)\,\rm{d}x-\int_{\Omega}\frac{(w_{1}-w_{2})^{2}}{2}(t_{2},\cdot)\,\rm{d}x+$ $\int_{t_{1}}^{t_{2}}\int_{\Omega}(q_{1}w_{1}-q_{2}w_{2})(w_{1}-w_{2})\,\rm{d}x\,\rm{d}t$
$=$ $-\int_{t_{1}}^{t_{2}}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}(w_{1}-w_{2})\right|^{2}\,\rm{d}x\,\rm{d}t=0.$
By Hardy-Littlewood-Sobolev estimate:
$\|(w_{1}-w_{2})\|_{\mathrm{L}^{2}(\Omega)}\leq C_{Vol}\|(w_{1}-w_{2})\|_{\mathrm{L}^{\frac{2n}{n-a}}(\Omega)}\leq C_{n,a,\Omega}\int_{\mathbb{R}^{N}}\left|(-\Delta)^{\frac{a}{2}}(w_{1}-w_{2})\right|^{2}\,\rm{d}x,$
where $C_{Vol}$ and $C_{n,a,\Omega}$ are two fixed positive constants depends on the domain and the dimension only. Hence
$\|w_{1}-w_{2}\|_{\mathrm{L}^{2}((t_{1},t_{2})\times\Omega)}=0,\quad\forall\,t_{1}\leq t_{2}\in[0,T].$
Hence
$w_{1}\equiv w_{2}.$
It further implies $(q_{1}-q_{2})w_{1}\equiv 0$. Let’s fix a time $t\in[0,T]$. We have
$\int_{\Omega}(q_{1}-q_{2})(t,\cdot)w_{1}(t,\cdot)\,\rm{d}x=0.$
Thanks to theorem 2, we have that $w_{1}(t,\cdot)$ is dense in $\mathrm{L}^{2}(\Omega)$. Hence $q_{1}(t,\cdot)\equiv q_{2}(t,\cdot)$. Since $t$ is arbitrary, we obtain
$q_{1}\equiv q_{2}.$
4. Further result:
In this section, we focus on a specific non-local elliptic eigenvalue Calderón problem and solve it using Pohozaev’s identity, as illustrated in [ROS14b]. We aim to demonstrate to the reader the significance of Pohozaev’s identity in solving various non-local Calderón problems. The problem discussed in this section is of an elliptic nature. We consider the eigenvalue problem
(34) {$(-\Delta)^{a}\phi=&\lambda\phi\qquad\qquad\mbox{in}\ (0,T)\times\Omega\\ \phi=&0\qquad\qquad\mbox{in}\ (0,T)\times\mathbb{R}^{N}\setminus\Omega.\end{aligned}\right.$
Here, the domain $\Omega\subset\mathbb{R}^{N}$ is a bounded $C^{1,1}$ domain with boundary $\partial\Omega$ and $a\in(0,1)$.
Theorem 10.
Let $\phi$ is a solution to the eigenvalue problem (34). Then for any non-empty smooth open set $\Sigma\subset\partial\Omega$
$\frac{\phi}{d^{a}}\Big{|}_{\Sigma}\neq 0.$
Before begin the proof, we begin with some available results. The set containing eigenvalues
[…]
is countable and discrete and the eigenvalues can be described in the following monotonic manner
$0<\lambda_{1}<\lambda_{2}\leq\lambda_{3}\leq\cdots\leq\lambda_{n}\leq\cdots+\infty.$
The following regularity of the eigenfunctions can be found in [ROS14a], [FRRO16]:
  • $\bullet$
    $\phi\in H^{a}(\mathbb{R}^{N})\cap C^{a}(\mathbb{R}^{N})$ and, for every $\beta\in[a,1+2a)$, $\phi$ is of class $C^{\beta}(\Omega)$ and
    (35) $[\phi(t,\cdot)]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
  • $\bullet$
    The function ${\frac{\phi}{d^{a}}}\Big{|}_{\Omega}$ can be continuously extended to $\overline{\Omega}$. Moreover, there exists $\alpha\in(0,1)$ such that ${\frac{\phi}{d^{a}}\in C^{\alpha}(\overline{\Omega})}$. In addition, for $\beta\in[\alpha,a+\alpha]$, the following estimate holds
    (36) $\left[\frac{\phi}{d^{a}}\right]_{C^{\beta}(\Omega_{\delta})}\leq C\delta^{a-\beta},\forall\delta\in(0,1).$
Hence Pohozaev identity (12) holds for $\phi$ [ROS14b].
(37) $\int_{\Omega}(x\cdot\nabla\phi)(-\Delta)^{a}\phi\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}$ $\phi(-\Delta)^{a}\phi\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{\phi}{d^{a}}\right)^{2}(x\cdot\nu)\,\rm{d}\sigma,$
where $\nu$ is the unit outward normal to $\partial\Omega$ at the point $x$ and $\Gamma$ is the gamma function. We state another version of Pohozaev identity. Let $\theta$ be another function such that $(-\Delta)^{a}\theta$ defined pointwise and ${\frac{\theta}{d^{a}}}$ continuously extended to boundary. Furthermore, let $\theta$ satisfy (35) and (36), then
(38) $\int_{\Omega}\theta_{x_{i}}(-\Delta)^{a}\phi\,\rm{d}x=-\int_{\Omega}\phi_{x_{i}}$ $(-\Delta)^{a}\theta\,\rm{d}x$
-- $\Gamma(1+a)^{2}\int_{\partial\Omega}\frac{\phi}{d^{a}}\frac{\theta}{d^{a}}\nu_{i}\,\rm{d}\sigma.$
Here the subscript $x_{i}$ denotes the derivative of the function with respect to the $i$th coordinate and $\nu_{i}$ is the $i$th component of the unit normal vector $\nu$ at the boundary. Proof of this can be found in [ROS14b].
Let $\alpha=(\alpha_{1},\cdot,\alpha_{N})\in\mathbb{R}^{N}$. For $1\leq i\leq N$, the function $\theta=\alpha_{i}\phi$ is an admissible function in (38). It yields
$\alpha_{i}\int_{\Omega}\phi_{x_{i}}(-\Delta)^{a}\phi\,\rm{d}x=-\alpha_{i}\int_{\Omega}\phi_{x_{i}}$ $(-\Delta)^{a}\phi\,\rm{d}x$
-- $\Gamma(1+a)^{2}\int_{\partial\Omega}\left(\frac{\phi}{d^{a}}\right)^{2}(\alpha_{i}\nu_{i})\,\rm{d}\sigma.$
It implies
$\alpha_{i}\int_{\Omega}\phi_{x_{i}}(-\Delta)^{a}\phi\,\rm{d}x=-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{\phi}{d^{a}}\right)^{2}(\alpha_{i}\nu_{i})\,\rm{d}\sigma,\quad\forall\,i=1,\cdots,N.$
Summing over the index ‘$i$’ yields
$\int_{\Omega}(\alpha\cdot\nabla\phi)(-\Delta)^{a}\phi=-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{\phi}{d^{a}}\right)^{2}(\alpha\cdot\nu)\,\rm{d}\sigma.$
Subtracting from (37), we obtain
(39) $\int_{\Omega}((x-\alpha)\cdot\nabla\phi)$ $(-\Delta)^{a}\phi\,\rm{d}x$
$=$ $\frac{2a-n}{2}\int_{\Omega}\phi(-\Delta)^{a}\phi\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{\phi}{d^{a}}\right)^{2}((x-\alpha)\cdot\nu)\,\rm{d}\sigma,\quad\forall\,\alpha\in\mathbb{R}^{N}.$
Next we move onto the proof of theorem 10.
Proof.
Let $\Omega\subset\mathbb{R}^{N}$ be a bounded $C^{1,1}$ domain with $\partial\Omega$ as it’s boundary. Let $\Sigma\subset\partial\Omega$ be a non-empty open set. Then there exists $\alpha\in\mathbb{R}^{N}$, such that [Uhl14]
$\Sigma:=$$\left\{x\in\partial\Omega,(x-\alpha)\cdot\nu\geq 0\right\}$
$\partial\Omega\setminus\Sigma:=$$\left\{x\in\partial\Omega,(x-\alpha)\cdot\nu\leq 0\right\}.$
We use the relation (39):
$\int_{\Omega}((x-\alpha)\cdot\nabla\phi)$ $(-\Delta)^{a}\phi\,\rm{d}x$
$=$ $\frac{2a-n}{2}\int_{\Omega}\phi(-\Delta)^{a}\phi\,\rm{d}x-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega}\left(\frac{\phi}{d^{a}}\right)^{2}((x-\alpha)\cdot\nu)\,\rm{d}\sigma.$
Plugging $(-\Delta)^{a}\phi=\lambda\phi$ in the above relation, we obtain
$\int_{\Omega}$ $\lambda(x-\alpha)\cdot\nabla\left(\frac{\phi^{2}}{2}\right)\,\rm{d}x=\frac{2a-n}{2}\int_{\Omega}\lambda\phi^{2}\,\rm{d}x$
$-\frac{\Gamma(1+a)^{2}}{2}\int_{\Sigma}\left(\frac{\phi}{d^{a}}\right)^{2}((x-\alpha)\cdot\nu)\,\rm{d}\sigma-\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega\setminus\Sigma}\left(\frac{\phi}{d^{a}}\right)^{2}((x-\alpha)\cdot\nu)\,\rm{d}\sigma.$
We prove via contradiction. Let ${\frac{\phi}{d^{a}}\Big{|}_{\Sigma}=0}$. Employing Gauss divergence theorem on the first term and thanks to ${\phi\equiv 0}$ outside $\Omega$, we obtain the following relation.
$a\lambda\int_{\Omega}\phi^{2}\,\rm{d}x=\frac{\Gamma(1+a)^{2}}{2}\int_{\partial\Omega\setminus\Sigma}\left(\frac{\phi}{d^{a}}\right)^{2}((x-\alpha)\cdot\nu)\,\rm{d}\sigma\leq 0$
As $a,\lambda>0$, it yields $\phi\equiv 0$, which contradicts $\lambda$ is an eigenvalue. Hence for any open subset $\Sigma\subset\partial\Omega$
$\frac{\phi}{d^{a}}\Big{|}_{\Sigma}\neq 0.$
While we were uploading our work to arXiv, Umberto Biccari, Mahamadi Warma, and Enrique Zuazua also posted their paper titled "Boundary Observation and Control for Fractional Heat and Wave Equations." In their work, they require the fractional power ${a\in[\frac{1}{2},1]}$ to achieve null controllability.
Acknowledgment: The author sincerely expresses gratitude to Prof. Tuhin Ghosh (Harish-Chandra Research Institute, Prayagraj, Uttar Pradesh, India) for his valuable insights and feedback, which significantly contributed to improving this article.